# Negative Sampling in Recommendation: A Survey and Future Directions

### v1

@@ 3.2. Dynamic Negative Sampling
With the aim to capture the relative informative negative samples in a universal manner, a wave of Dynamic Negative Sampling (DNS) strategies has sprung up in recent years. DNS refers to selecting the item that is more relevant to the positive sample representations (or user representations) as the negative samples from the dynamically selected item candidates. For example, the pioneering dynamic negative sampling strategy (Zhang et al., 2013) generally selects the hardest item from the randomly selected item candidates. The subsequent work primarily focuses on modeling the relationship between users and items, or precisely tuning the sampling difficulty of the candidate set. Particularly, DNS have outperformed most of the existing negative sampling methods in different recommendation objectives and scenarios.
@@
We categorize existing DNS strategies into six groups: Universal DNS, User-similarity DNS, Knowledge-aware DNS, Distribution-based DNS, Interpolation DNS and Mixed DNS. The comprehensive examination of their points of emphasis, advantages, and challenges is detailed in Table 3. Among these methodologies, Universal DNS stands out as an easy-to-deploy and universal method. However, its reliance on user-item matching scores poses a challenge to accuracy. User-similarity DNS prioritizes identifying user similarity associations and capturing user behavior similarities, while facing challenges in user association dependency, particularly for new users. Knowledge-aware DNS emphasizes samples with similar attributes to positives to capture their content correlations, introducing an additional dependency on knowledge. Distribution-based DNS analyzes the inherent distribution in the dataset, focusing on real negatives but adding the dependency on space complexity. Interpolation DNS balances positive and negative samples by injecting information from positives but faces the over-smoothing issue, neglecting crucial samples. Mixed DNS has exhibited significant flexibility by combining multiple strategies, but the complex hyperparameter tuning and substantial computational costs bring new challenges to these strategies. In the subsequent parts, we will introduce the implementation details and the effectiveness mechanisms of the existing dynamic negative sampling strategies for detecting informative negative samples to facilitate model optimization with enhanced accuracy and efficiency.
3.2.1. Universal DNS
Top-ranked universal negative selection. DNS (Zhang et al., 2013) hypothesizes that any unobserved item should not be ranked higher than any observed positive item and subsequently oversamples items top-ranked by the recommender from the randomly selected candidates. Due to the fact that the more popular an item is, the more times it acts as a positive sample in the conventional pairwise training procedure. DNS is expected to alleviate the popularity bias in RS. Along this line, a series of studies (Hu et al., 2018; Gong et al., 2022) enhance DNS from different perspectives. For example, MCRec (Hu et al., 2018) pretrains the nodes’ embedding with a pure matrix factorization method, computes and averages the pairwise similarities between two consecutive nodes along a path instance, and then keeps top-K path instances with the highest average similarities. To efficiently identify hard negative instances, CA-FME (Li et al., 2020) samples negative instances from triple sides of the positive triplet, and grants larger sampling probability for harder negative samples according to their compatibility scores.
Dynamic negative selection with high-ranking scope. PinSage (Ying et al., 2018) argues that the highest-scored item is likely to be a potentially positive sample, leading to the false negative problem. Meanwhile, it also demonstrates that the top-ranked items are more similar to the positive items than randomly selected ones, which forces the recommender to learn to distinguish items at a finer granularity. Specifically, it utilizes the query items’ PageRank score to rank the uniformly selected candidates, randomly samples hard negative items from the items ranked between 2000 and 5000, and designs a curriculum training scheme to help with convergence. ITSMGCN (Gong et al., 2022) proposes an improved dynamic negative sampler to detect the best sampling scheme by carefully tuning the number of item candidates and the ranking position of the hard negative sample in the ranking diagram. DNS (Shi et al., 2023) further reduces the probability of selecting overly hard negative samples by expanding both the number of item candidates and the selection range in DNS.
Collaborative distillation of dynamic negative samples. To address the sparsity of positive feedback and the ambiguity of missing feedback in recommendation, CD-SG (Lee et al., 2019) transposes DNS (Zhang et al., 2013) into knowledge distillation (KD) and proposes collaborative distillation (CD). It demonstrates that the higher-ranked items are more important because they can be potential inclusions in top-N recommendation. Specifically, the student model of CD-SG samples items with the dynamic soft target (the higher the ranking, the more the items are sampled) according to their ranking orders in the teacher model, with the aim to capture both positive and negative correlations and overcome the disadvantage of rank distillation that ignores negative feedback among items. Similarly, DE-RRD (Kang et al., 2020) reformulates the daunting task of learning all the precise ranking orders in recommendation to a relaxed ranking matching problem within the framework of KD. It first defines both the detailed ranking orders among the interesting items and the relative ranking orders between the interesting items and the uninteresting items from the teacher’s recommendation list and then defines a relaxed permutation probability to locate all the interesting items higher than all the uninteresting items in student model, while maintaining the detailed ranking orders among the interesting items.
Adjacent relation-based dynamic negative selection in graph structure. Along the development of GNN-based recommender models, a series of studies incorporated graph structure into the sampling strategy. For example, HIN-Bilinear (Hao et al., 2021b) leverages relative similar parts of the graph, constructs the sampled mini-batch with the sampled triplets and iteratively selects related items (users) of users (items) in sampled triplets to avoid ruining the topological structure of the large-scale graph. MCNS (Yang et al., 2020a) argues that the Markov chain for one node is likely to still work well for its neighbors due to the connatural property. It proposes an adapted version of the Metropolis-Hastings algorithm to traverse the graph structure by Depth First Search and generate negative samples from the last node’s Markov chain.
Dynamic negative selection with representation learning. With the advancement of representation learning in recommendation, a cohort of researchers and industry developers have embarked on an investigation of the distinction between hard negative samples and universal negative samples from the perspective of representation learning. To illustrate, PDRec (Ma et al., 2024a) examines the inherent information disparities among observed instances (positive samples) and the latent relevance of soft samples within the unobserved items (negative samples). It believes that following the emphasis on information-rich samples within the corpus as the positive instances, the deployment of hard negative sampling may concurrently exacerbate issues within the recommender, including data imbalance and over-fitting risks. To address this issue, PDRec designs a noise-free negative sampling strategy to capture items with low confidences as the Safe Negative Samples (SNS) which exhibit no relevance to user preferences in model optimization. Conversely, RealHNS (Ma et al., 2023) meticulously investigates the definition of Real Hard Negative Samples (RHNS) (samples that are neither too easy nor too hard) and False Hard Negative Samples (FHNS) (samples that represents users’ actual preferences in the future) and then defines the equation: HNS = RHNS + FHNS. To effectively leverage these two concepts, RealHNS employs a curriculum learning framework to construct the multi-granular RHNS selector (coarse-grained candidates selector and fine-grained samples selector). This facilitates the dynamic sampling of high-quality RHNS that are suitable for recommender’s current state, thereby alleviating the issue of negative transfer.
3.2.2. User-similarity DNS
User-similarity DNS identifies similar users based on their preferences from their historical behaviors and then dynamically selects items as negative samples according to this similarity association (Wu et al., 2019b; Wang et al., 2021b; Giobergia, 2022; Chen et al., 2019b). It can capture the dynamic user interest with historical user behaviors and approximate the user’s conditional preference state with the correlation between these users. Some studies emphasize the explicit associations among users, inferring the information dissemination path and collaborative relationships through the connections in social networks or explicit interaction behaviors. SamWalker (Chen et al., 2019b) proposes an efficient social-based sampling strategy to both speed up the gradient estimation and reduce the time-consuming of the preference inferring from the massive unobserved feedback. To reduce the extent to which the model relies on other users’ preferences, EvalRS (Giobergia, 2022) proposes a relatively audacious negative sampling strategy to capture neighboring users with their collaborative representations and regards item from the interaction list of the user’s neighbors that the user have not interacted as the negative sample. Conversely, other studies focus on implicit relationships, which refer to connections that are not directly observable. Such approaches aim to reveal the potential collaborative behaviors, interest similarities, and latent commonalities by mining users’ behavioral patterns. DEEMS (Wu et al., 2019b) believes that unobserved positive samples are more likely to exist correlations with observed positive samples than unobserved negative ones. Based on this hypothesis, DEEMS designs a hedging gradient between the user-centered and item-centered correlations to reorder the sampled unobserved items and rank the unobserved positive samples before the negative ones. To handle the absence of explicit social information, SamWalker++ (Wang et al., 2021b) constructs the pseudo-social network, that is, similar users are connected with specific item nodes or community nodes. It proposes a subtle sampling strategy based on the random walk along the pseudo-social network to precisely model the triplet confidence and determine which data are used to update parameters.
3.2.3. Knowledge-aware DNS
Knowledge-aware DNS dynamically captures the informative samples with their attributes or other relevant features and selects the informative items related to the positive sample as negative samples (Lian et al., 2020b; Chen et al., 2023b, 2016; Wang et al., 2021d). These strategies are designed to incorporate task-specific information to intelligently select more relevant and informative negative examples for the specific task. In Location Recommendation, users’ negatively-preferred locations are mixed with potentially positive ones in un-visited locations. However, informativeness differs from sample to sample, so treating them equally in optimization is far from optimal. GeoSAN (Lian et al., 2020b) designs geography-aware negative sampler to retrieve some nearest locations to the target location as the location candidates and employ negative sampling based on the uniform distribution or a popularity-based distribution. CITING (Chen et al., 2016) proposes a time-aware negative sampling strategy to select the tweets that were posted in close temporal proximity to his/her positive tweets as the negative samples in Tweet Recommendation. RPRM (Wang et al., 2021d) also designs a novel negative sampling strategy to model the agreement of review properties between the users’ interacted items and the unseen items. To investigate the Debiased Recommendation, FairNeg (Chen et al., 2023b) dynamically perceives the group-level unfairness based on the performance disparity of different category groups during the training process and adjusts each group’s corresponding negative sampling probability, intending to equalize all groups’ performance.
3.2.4. Distribution-based DNS
Distribution-based DNS analyzes the distribution pattern of positive and negative samples, then dynamically selects informative negative samples that do not overly disturb the training phase (Zhu et al., 2022; Ding et al., 2020; Lai et al., 2023). Some sampling strategies (Ding et al., 2020; Zhu et al., 2022) explore the statistical properties of both RHNS and FHNS to distinguish the relevant negative samples that are less informative or misleading, which are more likely to provide valuable information in RS. SRNS (Ding et al., 2020) discovers an empirical conclusion that only a limited number of samples are potentially crucial for the recommender, and false negative samples tend to get stable scores during training and proposes a variance-based sampling function to distinguish negative samples based on the aforementioned findings. GDNS (Zhu et al., 2022) then introduces an expectational gain-tuning sampler to leverage the high-gap expectation of users’ preference in the training process to reliably measure the quality of negative samples and dynamically guide the negative sampling procedure. In contrast, DENS (Lai et al., 2023) argues that real-world interactions are primarily driven by certain relevant factors associated with the items while not encompassing all factors. Therefore, DENS delineates the relevant and irrelevant factors within items that are associated with users or not. Subsequently, it devises a factor-aware sampling strategy to sample the appropriate by comparing the relevant factors while ensuring the similarity of irrelevant factors.
3.2.5. Interpolation DNS
Interpolation DNS synthesizes the informative negative samples by injecting information from positive samples with the interpolation technology (Huang et al., 2021; Shi et al., 2022). These approaches aim to explore the balance between maintaining the informative knowledge while avoiding excessive disturbance to the distribution of the feature space during the optimization process, which allows recommender to adapt to the changing requirements and optimize the balance between positive and negative samples. MixGCF (Huang et al., 2021) interpolates different proportions of positive embeddings into item candidates, optimizing the representation of the selected samples and translating the potential FNS into negative samples to avoid training bias. SDNS (Shi et al., 2022) proves that the positive-dominated mixing in MixGCF has the same effect as scaling the sigmoid function in BPR loss function in DNS, and specifically proposes an improvement scheme to define a soft BPR loss function, which not only preserves the integrity of the DNS (Zhang et al., 2013), but also yields remarkable improvements in terms of effectiveness and robustness.
3.2.6. Mixed DNS
Mixed DNS combines multiple dynamic negative sampling strategies to generate more diverse and effective negative sample candidates (Chen et al., 2017; Lei et al., 2020; Hidasi and Karatzoglou, 2018; Hidasi et al., 2015; Huang et al., 2020; Yang et al., 2020c; Xu et al., 2021). By integrating the uniform negative examples with those above dynamic negative sampling strategies, the existing Mixed DNS strategy provides a flexible framework that allows recommenders to customize the task-specific combination of different dynamic negative sampling strategies and attempt to overcome these limitations and achieve a more comprehensive and robust negative sampling process. As the pioneer Mixed DNS strategy, NNCF (Chen et al., 2017) proposes three novel sampling strategies to significantly improve training efficiency as well as the recommendation performance, named stratified sampling, negative sharing, and stratified sampling with negative sharing. Stratified sampling selects stratum (set of links on the bipartite graph sharing the same source or destination node) based on either item or user. Negative sharing constructs negative pairs by connecting non-linked users and items in the batch without any increasing computational burdens. And stratified sampling with negative sharing denotes the combination of the above two sampling strategies. For Conversational Recommendation, EAR (Lei et al., 2020) designs a mixed negative sampling strategy to simultaneously learn the user’s general and specific preference with two types of negative samples (the traditional randomly selected samples and the candidate samples satisfying the partially known preference in the conversation). FB-MNS (Huang et al., 2020) discovers that models trained with hard negative samples cannot outperform random negative samples. Therefore, It proposes a mixed sampling strategy to blend random and hard negatives and define the ratio of random to hard negative samples as in the training process, where the hard negative samples are selected between the rank position 101-500 to achieve the best Recall metric. GRU4Rec (Hidasi and Karatzoglou, 2018) extends the in-batch negative sampling strategy of GRU4Rec (Hidasi et al., 2015) with the additional shared samples in the mini-batch, where the shared samples are sampled with a popularity-based SNS strategy. MNS (Yang et al., 2020c) employs the same mixed sampling strategy as GRU4Rec (Hidasi and Karatzoglou, 2018) that incorporates the randomly selected additional samples as the supplement of the in-batch negative sampling strategy to tackle the selection bias and the inflexibility to adjust sampling distribution. Taking into account user feedback on attributes, FPAN (Xu et al., 2021) employs DNS (Zhang et al., 2013) to select informative negative samples in addition to directly sampling the non-interacted items as negative items, accelerating the training process.

### v2

@@ 3.2. Dynamic Negative Sampling
To capture the informative negative samples in a universal manner, a series of Dynamic Negative Sampling (DNS) strategies has sprung up in recent years. DNS refers to selecting the item that is more relevant to the positive sample representations (or user representations) as the negative samples from the dynamically selected item candidates. For example, the pioneering dynamic negative sampling strategy (Zhang et al., 2013) generally selects the hardest item from the randomly selected item candidates. The subsequent work primarily focuses on modeling the relationship between users and items, or precisely tuning the sampling difficulty of the candidate set. Particularly, DNS have outperformed most of the existing negative sampling methods in different recommendation objectives and scenarios. We categorize existing DNS strategies into six groups: Universal DNS, User-similarity DNS, Knowledge-aware DNS, Distribution-based DNS, Interpolation DNS and Mixed DNS. In the subsequent parts, we will introduce their implementation details and effectiveness mechanisms in detecting informative negative samples and facilitating recommender optimization.
3.2.1. Universal DNS
Top-ranked universal negative selection. DNS (Zhang et al., 2013) hypothesizes that any unobserved item should not be ranked higher than any observed positive item and subsequently oversamples items top-ranked by the recommender from the randomly selected candidates. Due to the fact that the more popular an item is, the more times it acts as a positive sample in the conventional pairwise training procedure. DNS is expected to alleviate the popularity bias in RS. Along this line, a series of studies (Hu et al., 2018; Gong et al., 2022) enhance DNS from different perspectives. For example, MCRec (Hu et al., 2018) pretrains the nodes’ embedding with a pure matrix factorization method, computes the pairwise similarities between two consecutive nodes along a path instance, and then keeps top-K path instances with the highest average similarities. To efficiently identify hard negative instances, CA-FME (Li et al., 2020) samples negative instances from triple sides of the positive triplet, and grants larger sampling probability for harder negative samples according to their compatibility scores.
Dynamic negative selection with high-ranking scope. PinSage (Ying et al., 2018) argues that the highest-scored item is likely to be a potentially positive sample, leading to the false negative problem. Meanwhile, it also demonstrates that top-ranked items are more similar to positive items than randomly selected ones, which forces the recommender to learn to distinguish items at a finer granularity. Specifically, it utilizes the query items’ PageRank score to rank the uniformly selected candidates, randomly samples hard negative items from the items ranked between 2000 and 5000. ITSMGCN (Gong et al., 2022) proposes an negative sampler to detect the best sampling scheme by carefully tuning the number of item candidates and the ranking position of hard negatives in ranking diagram. DNS (Shi et al., 2023) further reduces the probability of selecting overly hard negative samples by expanding both the number of item candidates and the selection range in DNS.
Collaborative distillation of dynamic negative samples. To address the sparsity of positive feedback and the ambiguity of missing feedback in recommendation, CD-SG (Lee et al., 2019) follows the knowledge distillation (KD) paradigm and proposes collaborative distillation (CD). Specifically, the student model of CD-SG samples items with the dynamic soft target according to their ranking orders in the teacher model, with the aim to capture both positive and negative correlations and overcome the disadvantage of rank distillation that ignores negative feedback among items. Similarly, DE-RRD (Kang et al., 2020) reformulates the task of learning all the precise ranking orders in recommendation to a relaxed ranking matching problem within the framework of KD. It first defines both the detailed ranking orders among the interesting items and the relative ranking orders between the interesting items and the uninteresting items from the teacher’s recommendation list and then defines a relaxed permutation probability to locate all the interesting items higher than all the uninteresting items in student model, while maintaining the detailed ranking orders among the interesting items.
Dynamic negative selection with representation learning. With the advancement of representation learning in recommendation, a cohort of researchers have embarked on an investigation of the distinction between easy and hard negative samples from the perspective of representation learning. To illustrate, PDRec (Ma et al., 2024a) examines the inherent information disparities among observed instances (positive samples) and the latent relevance within the unobserved items (negative samples). It believes that following the emphasis on information-rich samples within the corpus as the positive instances, the deployment of hard negative sampling may concurrently exacerbate the data imbalance and over-fitting risks within recommenders. To address this issue, PDRec designs a noise-free negative sampling strategy to capture items with low confidences as the Safe Negative Samples (SNS) which exhibit no relevance to user preferences in model optimization. In contrast, RealHNS (Ma et al., 2023) meticulously investigates the definition of Real Hard Negative Samples (RHNS) (samples that are neither too easy nor too hard) and False Hard Negative Samples (FHNS) (samples that represents users’ actual preferences in the future) and then defines the equation: HNS=RHNS+FHNS. To effectively leverage these two concepts, RealHNS employs a curriculum learning framework to construct the multi-granular RHNS selector (coarse-grained candidates selector and fine-grained samples selector). This facilitates the dynamic sampling of high-quality RHNS that are suitable for recommender’s current state, thereby alleviating the issue of negative transfer.
3.2.2. User-similarity DNS
User-similarity DNS identifies similar users based on their historical behaviors and then dynamically selects negative samples according to this similarity association (Wu et al., 2019b; Wang et al., 2021b; Giobergia, 2022; Chen et al., 2019b). Some studies emphasize the explicit associations among users to approximate users’ conditional preference state, inferring the information dissemination path and collaborative relationships through the connections in social networks or explicit interaction behaviors. SamWalker (Chen et al., 2019b) proposes an social-based sampling strategy to both speed up the gradient estimation and reduce the time-consuming of the preference inferring from the massive unobserved feedback. To reduce the dependency on other users’ preferences, EvalRS (Giobergia, 2022) proposes a relatively audacious negative sampling strategy to capture neighboring users with their collaborative representations and regards item from the interaction list of the user’s neighbors that the user have not interacted as the negative sample. Conversely, other studies focus on implicit relationships, which refer to connections that are not directly observable. Such approaches aim to reveal the potential collaborative behaviors and interest similarities by mining users’ behavioral patterns. DEEMS (Wu et al., 2019b) believes that unobserved positive samples are more likely to exist correlations with observed positives than unobserved negatives. Based on this hypothesis, DEEMS designs a hedging gradient between the user- and item-centered correlations to rerank the unobserved positive samples before the negative ones. To handle the absence of explicit social information, SamWalker++ (Wang et al., 2021b) constructs the pseudo-social network, that is, similar users are connected with specific item nodes or community nodes. It proposes a subtle sampling strategy based on the random walk along the pseudo-social network to precisely model the triplet confidence and determine which data are used to update parameters.
3.2.3. Knowledge-aware DNS
Knowledge-aware DNS dynamically captures the informative samples with their attributes or other relevant features and selects the informative items related to the positive sample as negative samples (Lian et al., 2020b; Chen et al., 2023b, 2016; Wang et al., 2021d). This strategy is designed to incorporate task-specific information to intelligently select more relevant and informative negative examples for the specific task. In Location Recommendation, users’ negatively-preferred locations are typically mixed with potentially positive ones. However, informativeness differs from sample to sample, so treating them equally in optimization is far from optimal. GeoSAN (Lian et al., 2020b) designs geography-aware negative sampler to retrieve some nearest locations to the target location as candidates and employ negative sampling based on the uniform distribution or a popularity-based distribution. CITING (Chen et al., 2016) proposes a time-aware negative sampling strategy to select the tweets that were posted in close temporal proximity to his/her positive tweets as negative ones in Tweet Recommendation. RPRM (Wang et al., 2021d) also designs a novel negative sampling strategy to model the agreement of review properties between the users’ interacted items and the unseen items. To investigate the Debiased Recommendation, FairNeg (Chen et al., 2023b) dynamically perceives the group-level unfairness based on their performance disparity during the training process and adjusts each group’s corresponding negative sampling probability, intending to equalize all groups’ performance.
3.2.4. Distribution-based DNS
Distribution-based DNS analyzes the distribution pattern of positive and negative samples, then dynamically selects informative negative samples that do not overly disturb the training phase (Zhu et al., 2022; Ding et al., 2020; Lai et al., 2023). Some sampling strategies (Ding et al., 2020; Zhu et al., 2022) explore the statistical properties of both RHNS and FHNS to distinguish the relevant negative samples that are less informative or misleading. SRNS (Ding et al., 2020) discovers an empirical conclusion that only a limited number of samples are potentially crucial for the recommender, and false negative samples tend to get stable scores during training and further proposes a variance-based sampling function to distinguish negative samples. GDNS (Zhu et al., 2022) introduces an expectational gain-tuning sampler to leverage the high-gap expectation of users’ preference in the training process to measure the quality of negative samples and dynamically guide the negative sampling procedure. In contrast, DENS (Lai et al., 2023) argues that real-world interactions are primarily driven by certain relevant factors associated with the items while not encompassing all factors. Therefore, It delineates the relevant and irrelevant factors within items that are associated with users or not and then devises a factor-aware sampling strategy to sample the appropriate by comparing the relevant factors while ensuring the similarity of irrelevant factors.
3.2.5. Interpolation DNS
Interpolation DNS synthesizes the informative negative samples by injecting information from positive samples with the interpolation technology (Huang et al., 2021; Shi et al., 2022). These approaches aim to explore the balance between maintaining the informative knowledge while avoiding excessive disturbance to the distribution of the representation space during optimization, allowing recommender to adapt to the changing requirements and maintain the balance between positive and negative samples. MixGCF (Huang et al., 2021) interpolates different proportions of positive embeddings into item candidates, optimizing the representation of the selected samples and translating the potential FNS into negative samples to avoid training bias. SDNS (Shi et al., 2022) proves that the positive-dominated mixing in MixGCF has the same effect as scaling the sigmoid function in BPR loss function in DNS, and specifically proposes an improvement scheme to define a soft BPR loss function, which not only preserves the integrity of the DNS (Zhang et al., 2013), but also yields remarkable improvements in terms of effectiveness and robustness.
3.2.6. Mixed DNS
Mixed DNS combines multiple DNS strategies to generate more diverse and effective negative sample candidates (Chen et al., 2017; Lei et al., 2020; Hidasi and Karatzoglou, 2018; Hidasi et al., 2015; Huang et al., 2020; Yang et al., 2020c; Xu et al., 2021). By integrating the uniform negative examples with those above DNS strategies, the existing Mixed DNS strategy provides a flexible framework that allows recommenders to customize the task-specific combination of different DNS strategies and attempt to achieve a more comprehensive and robust negative sampling process. As the pioneering Mixed DNS strategy, NNCF (Chen et al., 2017) proposes three novel sampling strategies to significantly improve training efficiency as well as the recommendation performance, named stratified sampling, negative sharing, and stratified sampling with negative sharing. Stratified sampling selects stratum (set of links on the bipartite graph sharing the same source or destination node) based on either item or user. Negative sharing constructs negative pairs by connecting non-linked users and items in the batch without any increasing computational burdens. And stratified sampling with negative sharing denotes the combination of the above two sampling strategies. For Conversational Recommendation, EAR (Lei et al., 2020) designs a mixed negative sampling strategy to simultaneously learn the user’s general and specific preference with two types of negative samples (the traditional randomly selected samples and the samples satisfying the partially known preference in the conversation). FB-MNS (Huang et al., 2020) discovers that models trained with hard negative samples cannot outperform random negative samples. Therefore, It proposes a mixed sampling strategy to blend random and hard negatives with the ratio as in the training process, where the hard negative samples are selected between the rank position 101-500 to achieve the best Recall metric. GRU4Rec (Hidasi and Karatzoglou, 2018) extends the in-batch negative sampling strategy of GRU4Rec (Hidasi et al., 2015) with the additional shared samples in the mini-batch, where the shared samples are sampled with a popularity-based SNS strategy. MNS (Yang et al., 2020c) employs the same mixed sampling strategy as GRU4Rec (Hidasi and Karatzoglou, 2018) that incorporates the randomly selected additional samples as the supplement of the in-batch negative sampling strategy to tackle the selection bias and the inflexibility to adjust sampling distribution. Taking into account user feedback on attributes, FPAN (Xu et al., 2021) employs DNS (Zhang et al., 2013) to select informative negative samples in addition to directly sampling the non-interacted items as negative items, accelerating the training process.
@@ 3.2.6. Mixed DNS
Summary: The comprehensive examination of the emphasis, advantages, and challenges of six types of DNS algorithms is detailed in Table 3. Among these methodologies, Universal DNS stands out as an easy-to-deploy and universal method. However, its reliance on user-item matching scores poses a challenge to accuracy. User-similarity DNS prioritizes identifying user behavior similarities, while facing challenges in user association dependency, particularly for new users. Knowledge-aware DNS emphasizes samples with similar attributes to positives to capture their content correlations, introducing an additional dependency on knowledge. Distribution-based DNS analyzes the inherent distribution in the dataset, focusing on real negatives but adding the dependency on space complexity. Interpolation DNS balances positive and negative samples by injecting information from positives but faces the over-smoothing issue, neglecting crucial samples. Mixed DNS has exhibited significant flexibility by combining multiple strategies, but the complex hyperparameter tuning and substantial computational costs bring new challenges to these strategies.